UT 20130201 Technical Observing

Observer: JHill, MDeLaPena, CBiddick, DSummers (Tucson)
Telescope Operator: DHuerta (LBTO)
Additional AO Support: DMiller (LBTO)
Telescope Support: JLittle (LBTO) DAshy, DCox, TSargent (Tucson)
Instrument Support: JMorris (LBTO)

Summary

Cloudy at the beginning. Then problems with telescope out of balance. Then problems with rotators. Finally opened dome for LBC tests with 2013A at local midnight. Switched over to LBTI on left side only at 9UT (There are some advantages to not having a rotator.)

Details

Problems on top of problems

01:57 Mirror Ventilation turned on first time. (It was on and off several times between other activities.)

04:00 Sky has cleared.

04:30 Telescope is balanced.

06:40 Last of the instrument rotators declared dead. See details in Issue 4473.

06:45 Reconfiguring swing arms to LBCs.

LBC Tests with Build 2013A

Open

07:00 Open shutter doors

07:10 Mirror Ventilation turned on again - now with shutter doors open.

07:15 LBC BOTH TRACK preset to WT10_0301

07:20 replay preset x200 to get guiding Blue is 3.2 arcsec w/o collimation, Red is much worse (collimation LUT possibly not updated for new hub position).

07:25 DOFPIA (converges on Red from a poor starting condition)

07:32 preset x200 again (some hints of astigmatism / coma in blue) EL=55 The Blue pupils have some hint of a skirt.

Red says V=BESSEL (?)

07:37 Fast OB with 4 500 arcsec dithers and single 240 sec exposures. We are doing these large dithers to explore Issue 4450.

07:58 preset ACT0348 EL=78

08:01 DOFPIA

08:05 Fast OB with 4 500 arcsec dithers and single 240 sec exposures. (Interesting Drifts)

08:25 NSI preset to C/2009 P1 EL=44 dRA=-43 arcsec/hour

Did not run DOFPIA before the comet images.....

08:28 Fast OB with 3 dithers and single 240 sec exposures on the NS target. NS Guiding residuals look good. The comet nucleus is round while the stars are trailed.

Third exposure at ~8.66 UT has a 0.6 arcsec jump in EL showing on both Red and Blue. The big guide correction on Red was at 8.671 UT while Blue perhaps split two corrections near 8.67 UT.

Outside temperature saw a 4 degC temperature glitch (because of wind data) at 8.665 UT. I'm suspicious that this caused the jump. i've reopened Issue 4459.

LBC tests with Build 2013A have been successfully completed.

Close

08:44 To zenith for reconfiguring to Bent Gregorian (authorizing LBTI on Left only)

08:52 pair of LBC biases with dome lights on

LBTI on SX only

09:00 Parked at zenith with lights on for LUCI1 communication problem and LBTI alignment.

Open

10:40 ADAPTIVEACE_TRACK preset from LBTI @ LCBG

.....not taking notes here...working on HDF Telemetry plotting scripts.......

12:05 LBTI has a star image, but there is some problem communication to their left side WFS. We think that is why PresetAO is failing.

12:40 JMH departs

Close

13:45 Close Shutter Doors

-- JohnHill - 01 Feb 2013
Topic attachments
I Attachment Action Size Date Who Comment
daq.20130201telescope.jpgjpg daq.20130201telescope.jpg manage 703 K 05 Feb 2013 - 21:38 JohnHill Plot of Telescope Temperatures
guidelbc.sx.20130201lbto.jpgjpg guidelbc.sx.20130201lbto.jpg manage 334 K 06 Feb 2013 - 02:55 JohnHill Plot of LBC Guiding Corrections
Topic revision: r7 - 06 Feb 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback